# Efficient Action Counting with Dynamic Queries

### v2

@@ Abstract
Temporal repetition counting aims to quantify the repeated action cycles within a video. The majority of existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity. In this work, we introduce a novel approach that employs an action query representation to localize repeated action cycles with linear computational complexity. Based on this representation, we further develop two key components to tackle the essential challenges of temporal repetition counting. Firstly, to facilitate open-set action counting, we propose the dynamic update scheme on action queries. Unlike static action queries, this approach dynamically embeds video features into action queries, offering a more flexible and generalizable representation. Secondly, to distinguish between actions of interest and background noise actions, we incorporate inter-query contrastive learning to regularize the video representations corresponding to different action queries. As a result, our method significantly outperforms previous works, particularly in terms of long video sequences, unseen actions, and actions at various speeds. On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction. Code is available at https://github.com/lizishi/DeTRC.
@@
Temporal repetition counting Video understanding
@@ 1 Introduction
Recent progress in action detection [33, 23] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [3]. Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query. This novel formulation reduces the complexity from quadratic to linear and enables counting long videos with varying action periods, as illustrated in Fig. 1 (b). However, directly applying the action detection approaches [23, 44, 33] to the TRC problem proves inadequate (Tab. 1) in addressing two distinctive challenges unique to TRC. These challenges underscore the complexity of TRC, highlighting why TRC is not merely another action detection task but requires a nuanced approach that considers the unique nature of repetitive actions. We highlight the two inherent differences between TRC and the classical action detection task:
@@
Figure 1: Comparison of (a) similarity-matrix-based methods and (b) our proposed action query-based method. Unlike the similarity-matrix-based methods, which exhibit a quadratic growth $O(T^{2}C)$ in computational complexity as the length of the input video $T$ increases, our approach is characterized by a linear growth $O(TC)$ in complexity, where $C$ denotes the feature dimension of the video frames. This distinct advantage allows us to effectively process long videos that contain varying durations of action sequences.
As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec. 4.3. In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC. In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using distilled content features from the video encoder. This mechanism allows the decoder to attend to the “action of interest” based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition. To tackle the second challenge, we propose Inter-query Contrastive Learning (ICL). It enforces that the repetitive action cycles are grouped together in the learned representation space while being separated apart from non-repetitive (e.g. background or distracting) video content. The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity. In other words, our queries are designed to localize the contextually similar action instances, which aligns exactly with the definition of repetition counting. Extensive experiments validate the effectiveness of the two proposed designs.
@@
    We propose Dynamic Action Query to guide the model to focus on the action of interest and improve generalization ability across different actions.
@@
    We introduce Inter-query Contrastive Learning to facilitate learning repetitive action representations and to distinguish these from distractions.
@@ 3.1 Overview
Given an RGB video sequence $\mathbf{V}$ with $T$ frames, The TRC task aims to predict an integer $N$ indicating the number of estimated repetitive action cycles. Our model consists of a video backbone network $\Phi(\cdot)$, a Transformer network with encoder $\mathcal{E}(\cdot)$, decoder $\mathcal{D}(\cdot)$, and multiple prediction heads as shown in Fig. 2. The backbone network $\Phi(\cdot)$ takes a sequence of $T$ video frames as input and extracts feature vectors $\mathbf{F}\in\mathbb{R}^{T\times C}$ for each frame, where $C$ denotes the feature dimension. In practice, we utilize the commonly-used backbone works, e.g. TSN [40], and freeze the backbone parameters.
Figure 2: Framework overview. Given a video input, we first extract video features $\mathbf{F}$ with the backbone network Φ()normal-$\Phi(\cdot)$. Then we leverage a Transformer encoder-decoder network to process the features $\mathbf{F}$ to detect each action cycle instance by predicting its action class label $\mathbf{C}^{\text{cls}}$ and its time position $\mathbf{P}$. To tackle the challenge of open-set action categories, we propose Dynamic Action Query (DAQ), allowing the method to attend to the “action of interest” based on the input video contents in a dynamic and contextually aware manner. To recognize actions with identical content, we propose Inter-query Contrastive Learning (ICL) to group the “action of interest” and be separated apart from noise actions (e.g. background or distractions). Integrating DAQ and ICL, our method can identify contextually similar action instances that are adaptively conditioned on the video content, without manual action class definition. By counting the queries classified as the “action of interest” with classification confidence score $\mathbf{C}^{\text{conf}}$ exceeding the threshold $\alpha$, we finally get the total count value $N$.
We then design a Transformer encoder-decoder architecture to process the features $\mathbf{F}$ to detect each action cycle instance and obtain the final action cycle count value $N$. This requires the model to discern each detected action cycle class $\mathbf{C}^{\text{cls}}$ and its time position $\mathbf{P}$. To achieve this, we introduce the concept of action queries within the decoder $\mathcal{D}(\cdot)$, which are comprised of content queries $\mathbf{Q}^{\text{cnt}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, where $Q$ represents the number of queries. Refer to Fig. 2 for illustration. The content queries $\mathbf{Q}^{\text{cnt}}$ are responsible for capturing the action category, while the position queries $\mathbf{Q}^{\text{pos}}$ handle the action interval estimation. We use two prediction heads including a classification head and a position head after the decoder to predict the class label $\mathbf{C}^{\text{cls}}$ and time position $\mathbf{P}$ for each query, respectively.
To tackle the challenge of open-set action categories, we define $K=2$ classes, namely “action of interest” and “null ($\emptyset$)”. We then propose the Dynamic Action Query (DAQ) strategy which adaptively updates the content query $\mathbf{Q}^{\text{cnt}}$ of the decoder using distilled content features that are selected from the encoder $\mathcal{E}(\cdot)$. This mechanism eliminates the need for a manual definition of the specific semantics of “action of interest” and allows the decoder to attend to the “action of interest” conditioned on the input video contents in a dynamic and contextually-aware manner. Nevertheless, videos may contain distracting factors such as background elements or other noise actions, necessitating that the detected “action of interest” instances exhibit identical or similar contents. To this end, we propose Inter-query Contrastive Learning (ICL) to cluster content queries into positive action set ($S^{+}$) and negative null set ($S^{-}$) based on their predicted class label $\mathbf{C}^{\text{cls}}$. Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time. Finally, by counting the queries classified as the “action of interest” with classification confidence score $\mathbf{C}^{\text{conf}}$ exceeding a threshold $\alpha$, we get the total count value $N$.
@@ 3.2 Model Architecture
Encoder. The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [39] architecture which processes the video features $\mathbf{F}$ extracted by the backbone network. The output features $\mathbf{H}^{\text{enc}}$ of $\mathcal{E}(\cdot)$ will then be fed into a query selection module to select $Q$ queries to be the initial content queries $\mathbf{Q}^{\text{cnt}}$ for the decoder. For clarity, we will detail the query selection module later in this section. At now, it can be considered as a rough filter that selects top-$Q$ relevant features as the input queries for decoder $\mathcal{D}(\cdot)$.
Decoder. The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [39] architecture, which takes the encoder features $\mathbf{H}^{\text{enc}}$ and two types of action queries as input and outputs a processed feature $\mathbf{H}^{\text{dec}}\in\mathbb{R}^{Q\times C}$. The action queries are constructed by content queries $\mathbf{Q}^{\text{cnt}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, where $\mathbf{Q}^{\text{cnt}}$ and $\mathbf{Q}^{\text{pos}}$ are designed for representing the content features and the time position of each action cycle, respectively. The content queries $\mathbf{Q}^{\text{cnt}}$ are directly initialized from the distilled content features selected from the encoder $\mathcal{E}(\cdot)$, i.e. the proposed DAQ strategy, which will be detailed in Sec. 3.3. The position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters and transformed to the initial time position $\mathbf{P}^{\rm init}=(\mathbf{m}^{\rm init},\mathbf{d}^{\rm init})\in\mathbb{R}^{Q\times 2}$ with a linear layer before feeding into the decoder $\mathcal{D}(\cdot)$, where we propose to use the midpoint $\mathbf{m}$ and the period duration $\mathbf{d}$ to denote each action time position. For clarity, we omit the linear layer in Fig. 2.
Prediction heads. The prediction heads consist of a classification head network and a position head network, which are both MLPs.
The classification head takes the output features of the decoder $\mathbf{H}^{\text{dec}}$ as input and predicts the action class label for each content query. Given that the action category is an open set and not preset, we set the number of classes $K$ to be 2 in the classification head, indicating whether it is an “action of interest” or a “null ($\emptyset$)” class. The classification head employs the Softmax function to calculate the probabilities $\mathbf{C}^{\text{cls}}$ of $K$ categories for each action query. We then designate the category with the highest probability as the predicted class label for the respective action query. Subsequently, we regard this probability value as the estimated confidence score $\mathbf{C}^{\text{conf}}$. We define an action query instance as “action of interest” when its confidence score $\mathbf{C}^{\text{conf}}$ is higher than the threshold $\alpha$.
We apply binary Cross-entropy loss to supervise the predicted class labels $\mathbf{C}^{\text{cls}}$ for each action instance, aiming to guide the model in effectively discriminating the action instances of interest from others:
$\mathcal{L}_{\text{cls}}=\sum\limits_{q=1}^{Q}{\rm CrossEntropyLoss}\left(\mathbf{C}^{\text{cls}}_{q},\tilde{\mathbf{C}}^{\text{cls}}_{q}\right)=\sum\limits_{q=1}^{Q}\sum\limits_{i=1}^{K}-\tilde{c}_{q,i}{\rm log}\left(c_{q,i}\right),$ (1)
where $\mathbf{C}^{\text{cls}}_{q}=[c_{q,1},c_{q,2},...,c_{q,K}]\in\mathbb{R}^{K}$ represents the predicted class label for $q^{th}$ action query. $\tilde{}$ denotes the GT label.
The position head estimates the residual of the midpoint and duration based on the decoder features $\mathbf{H}^{\text{dec}}$ for each query, and adds to the initial time position $\mathbf{P}^{\rm init}$ to get the final estimated time position. We supervise the predicted time positions with $L_{1}$ loss and gIoU [31] loss:
@@
$\begin{split}\mathcal{L}_{\text{pos}}&=\sum\limits_{q=1}^{Q}\left\|\mathbf{P}_{q}-\tilde{\mathbf{P}}_{q}\right\|+\lambda_{\text{gIoU}}\sum\limits_{q=1}^{Q}\left(1-{\rm gIoU}(\mathbf{P}_{q},\tilde{\mathbf{P}}_{q})\right),\end{split}$ (2)
where $\mathbf{P}_{q}=(\mathbf{m}_{q},\mathbf{d}_{q})$ represents the predicted action instance position for the $q^{th}$ query. $\tilde{}$ stands for GT label.
Note that the $Q$ predicted instances from the prediction heads are in a random order. Therefore, before applying the two loss terms, we have to establish a one-to-one correspondence between the disordered predicted instances and the ground-truth (GT) labels. Concretely, we employ the Hungarian matching algorithm [15] and the matching process considers two criteria: 1) the action class label is the “action of interest”, and 2) the time position of the action must align with a corresponding ground-truth time position label. A matching cost matrix is constructed based on these criteria. For the first criterion, the Negative Log-Likelihood (NLL) loss is employed to assess classification accuracy. For the second aspect, both Intersection over Union (IoU) and $L_{1}$ distance are utilized to compute costs, measuring the alignment between the estimated time positions and GT labels. The matching cost matrix is of dimensions $Q\times N_{gt}$, where $N_{gt}$ represents the number of GT action instances. Utilizing the Hungarian matching algorithm [15] based on this cost matrix, we derive the prediction-to-GT matching results.
Query selection. As we mentioned above, videos may contain multiple action categories or other distracting factors such as background elements, therefore, we design a query selection module after the encoder $\mathcal{E}(\cdot)$ to filter the “action of interest” coarsely. To achieve this, we reuse the two prediction heads to predict the class labels and time positions for the output features $\mathbf{H}^{\text{enc}}$ of the encoder $\mathcal{E}(\cdot)$. Based on their predicted confidence scores $\mathbf{C}^{\text{conf}}$, the query selection module selects top-$Q$ queries with high confidence that are classified as the “action of interest” to be the initial content queries $\mathbf{Q}^{\text{cnt}}$ for the decoder. For clarity, we omit the encoder prediction heads in Fig. 2.
Note that the two prediction head networks are shared between the encoder and decoder. In addition, we apply corresponding supervision using $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{pos}}$ to the encoder prediction heads. Additional experimental results of the encoder prediction heads can be found in the supplementary materials.

### v3

@@ Abstract
Most existing methods rely on the similarity correlation matrix to characterize the repetitiveness of actions, but their scalability is hindered due to the quadratic computational complexity. In this work, we introduce a novel approach that employs an action query representation to localize repeated action cycles with linear computational complexity. Based on this representation, we develop two key components to tackle the essential challenges of temporal repetition counting. Firstly, to facilitate open-set action counting, we propose the dynamic update scheme on action queries. Unlike static action queries, this approach dynamically embeds video features into action queries, offering a more flexible and generalizable representation. Secondly, to distinguish between actions of interest and background noise actions, we incorporate inter-query contrastive learning to regularize the video representations corresponding to different action queries. As a result, our method significantly outperforms previous works, particularly in terms of long video sequences, unseen actions, and actions at various speeds. On the challenging RepCountA benchmark, we outperform the state-of-the-art method TransRAC by 26.5% in OBO accuracy, with a 22.7% mean error decrease and 94.1% computational burden reduction. Code will be publicly available.
@@
Temporal repetition counting, Video understanding
@@ 1 Introduction
Figure 1: Conceptual workflow comparison of (a) similarity-matrix-based methods and (b) our proposed action query-based method. (a) Most existing methods use similarity matrices calculated between each frame to detect repetitive actions, resulting in a time complexity of $O(T^{2}C)$, where $T$ denotes video length and $C$ denotes feature dimension. (b) On the other hand, our method employs action queries to represent each action cycle and estimates their classes and temporal locations. Each query is classified into either “repetitive actions” or “others ($\emptyset$)” class labels. The temporal location is expressed by the midpoint and duration on the timeline. During training, we perform bipartite matching to uniquely associate a prediction to a GT action cycle, taking into account both class labels and temporal locations. Predictions with no match should yield a $\emptyset$ class prediction, indicating that these are not repetitive actions. This novel formulation reduces the complexity from quadratic to linear as $O(TC)$.
@@
Recent progress in action detection [34, 24] introduces an efficient representation of action periods by associating each action instance with an action query, similar to DETR [3]. Inspired by this, we propose to formulate the TRC problem as a set prediction task where the goal is to detect every action cycle by representing it as an action query, as illustrated in Fig. 1 (b). Based on this query representation, we use a Transformer encoder-decoder network [50] to detect repetitive action instances and their temporal positions, defined by their midpoints and durations. This novel formulation reduces the complexity from quadratic to linear  [^1] In implementation, we employ deformable attention modules as proposed in DeformableDETR [50]. and enables counting long videos with varying action periods. However, directly applying the action detection approaches [24, 45, 34] to the TRC problem proves inadequate (Tab. 1) in addressing two distinctive challenges unique to TRC. These challenges underscore the complexity of TRC, highlighting why TRC is not merely another action detection task but requires a nuanced approach that considers the unique nature of repetitive actions. We highlight the two inherent differences between TRC and the classical action detection task:
@@
As a result, approaching TRC as a simple action detection task results in inferior performance as shown in Sec. 4.3. In contrast, we propose two novel strategies to address these challenges and redefine the framework for TRC. In response to the first challenge, we propose Dynamic Action Query (DAQ), which adaptively updates the action query using content features extracted from the video (Fig. 2). This mechanism allows the decoder to attend to the repetitive actions based on the input video contents in a dynamic and contextually aware manner, without the need for manual definition. To tackle the second challenge, we define two action classes: “repetitive actions”, which are dynamically defined by DAQ, and the “others ($\emptyset$)” class. We further propose Inter-query Contrastive Learning (ICL) which ensures that primary repetitive action cycles are grouped together in the learned representation space while being separated from other distractors, such as background noise. The integration of two core components (DAQ and ICL) ensures that the action instances are identified adaptively based on the video content and their contextual similarity. In other words, our queries are designed to localize the contextually similar action instances, which aligns exactly with the definition of repetition counting. Extensive experiments validate the effectiveness of the two proposed designs.
@@
    We propose Dynamic Action Query to guide the model to focus on the repetitive actions contextually defined by the video content thereby improving generalization ability across different actions.
@@
    We introduce Inter-query Contrastive Learning to facilitate learning primary repetitive action representations and to distinguish them from distractions.
@@ 3 Method
3.1 Preliminary
DETR [3] is a pioneering object detection framework that builds upon the Transformer encoder-decoder architecture [40]. The overall DETR architecture [3] consists of three main components: a backbone to extract image features, an encoder-decoder Transformer, and the detection heads, i.e. feed-forward network (FFN) that makes the final detection prediction. The main features of DETR are the conjunction of a bipartite matching loss and transformers with (non-autoregressive) parallel decoding. The bipartite matching loss is a set-based Hungarian loss during training, which uniquely assigns a prediction to a GT object, and is invariant to a permutation of predicted objects. This design enables DETR to perform parallel processing and predict all objects simultaneously. We briefly review the workflow as follows.
Given the input image feature maps extracted by a CNN backbone, e.g. ResNet [13], DETR exploits a standard Transformer encoder-decoder network to transform the feature maps to be features of a set of object queries. An FFN and a linear projection are added on top of the object query features as the detection heads. The FFN acts as the regression branch that predicts the bounding box coordinates, i.e. box center coordinates, box height and width. The linear projection acts as the classification branch to produce the classification results, i.e. $\mathtt{object}$ vs. $\mathtt{no\,object}\,(\emptyset)$. The $\emptyset$ class is used to represent that no object is detected, playing a similar role to the “background” class in the standard object detection approaches.
DETR infers a fixed-size set of $N$ predictions in a single pass, where $N$ is set to be significantly larger than the typical number of objects in an image. During training, a Hungarian loss produces an optimal bipartite matching between predicted and GT objects and then optimizes the object position-specific losses. The matching procedure takes into account both the class prediction and the similarity of predicted and GT boxes and finds one-to-one matching for direct set prediction without duplicates.
Figure 2: Overview and detailed architecture design of our method. (Left) We utilize a DETR-inspired framework to predict the number of repetitive action cycles in an RGB video sequence. The architecture consists of a backbone network $\Phi(\cdot)$ for feature extraction, an encoder-decoder Transformer for processing query features, and prediction heads for action classification. By counting the queries classified as “repetitive actions”, we get the total count $N$. (Right) The detailed workflow within the black box on the left. Given the video features $\mathbf{F}$ as input, the encoder $\mathcal{E}(\cdot)$ produces two sets of query features, i.e. action queries and position queries. A query selection module then screens these queries, retaining only the most relevant ones for forwarding to the decoder $\mathcal{D}(\cdot)$. This process utilizes action classification results from prediction heads positioned after the encoder. The decoder adopts the Dynamic Action Query (DAQ) strategy, utilizing the selected action queries as priors, while incorporating a learnable position query and attending to the encoder’s output. The output embeddings from the decoder are then passed to their corresponding prediction head, which estimates the class label $\mathbf{A}$ and temporal location $\mathbf{P}$. During training, we employ bipartite matching to uniquely pair each prediction with a GT action cycle $\hat{\mathbf{P}}$, considering both class labels and temporal locations. Predictions that fail to match should yield $\emptyset$ class, signifying that they do not represent repetitive actions. The DAQ strategy addresses the open-set problem by contextually defining “repetitive actions” based on video features. To further distinguish these actions from $\emptyset$ distractions, we propose Inter-query Contrastive Learning (ICL). By integrating DAQ and ICL, our method effectively identifies contextually similar actions and excludes distractions, making it well-suited for tackling the challenges of the TRC problem.
@@ 3.2 Overview
Given an RGB video sequence with $T$ frames, the TRC task aims to predict an integer $N$ indicating the number of detected repetitive action cycles. Drawing inspiration from DETR [3], we streamline the problem as a set detection task in the temporal domain. The overall framework consists of three main components: a backbone network $\Phi(\cdot)$, an encoder-decoder Transformer, and prediction heads, as shown in Fig. 2. The backbone extracts video features $\mathbf{F}$, which is transformed to features of a set of action queries by an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$. Two prediction heads are added on top of the query features. The prediction heads consist of an action head and a position head. The action head classifies the query, while the position head estimates its temporal location.
Building on previous enhancements in DETR models [23], we introduce two distinct types of queries to enhance the query-to-feature similarity and accelerate convergence: action queries and position queries. Action queries are designed to capture features pertinent to action classification, whereas position queries concentrate on the temporal dimensions of an action. As depicted in the detailed diagram on the right side of Fig. 2, two specialized prediction heads (i.e. action head and position head) are tasked with processing these respective sets of queries to generate predictions. To further accelerate convergence, we introduce a post-encoder query selection module that acts as an initial sieve, filtering out the majority of queries unlikely to represent repetitive actions before forwarding the most promising ones to the decoder. This enhances the efficiency and focus of the subsequent prediction process.
Considering the two challenges of the TRC tasks discussed in Sec. 1, we propose Dynamic Action Query (DAQ) strategy to address the open-set problem. DAQ directly uses dynamically updated action query features from the encoder to initialize the action query of the decoder. This strategy defines what “repetitive actions” should be detected based on the video content features in a dynamic and contextually aware manner. To tackle the second challenge, we define two action classes: “repetitive actions” and the “others ($\emptyset$)” class. We further propose Inter-query Contrastive Learning (ICL) to cluster repetitive action queries into a positive action set ($S^{+}$) and group the other queries to a negative set ($S^{-}$) in the feature space. Integrating DAQ and ICL allows our method to identify contextually similar action instances that are adaptively based on the video content, and exclude other distracting actions at the same time.
Similar to DETR [3], during training, we employ bipartite matching which uniquely assigns a prediction to a GT action cycle. The matching procedure takes into account both the class prediction and the similarity of predicted and GT temporal locations. We then optimize the action cycle-specific losses. In inference, our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be significantly larger than the typical number of action cycles in a $T$-frame video sequence. By counting the queries classified as “repetitive actions”, we get the final total count value $N$.
In the following, we will introduce the network architecture design in Sec. 3.3, the DAQ and ICL modules in Sec. 3.4 and Sec. 3.5, and the model training in Sec. 3.6.
@@ 3.3 Model Architecture
Backbone. The backbone network $\Phi(\cdot)$ takes a sequence of $T$ video frames as input and extracts feature vectors $\mathbf{F}\in\mathbb{R}^{T\times C}$ for each frame, where $C$ denotes the feature dimension.
Encoder. The encoder $\mathcal{E}(\cdot)$ is a classical Transformer [40] architecture which has ${L}^{\text{enc}}$ standard encoder layers. The encoder transforms the video features $\mathbf{F}$ into two distinct query types: action queries $\tilde{\mathbf{E}}^{\text{act}}\in\mathbb{R}^{T\times C}$ and position queries $\tilde{\mathbf{E}}^{\text{pos}}\in\mathbb{R}^{T\times C}$, as shown in Fig. 2 (right). This design enhances the query-to-feature similarity and accelerates convergence, where action queries capture features pertinent to action classification. On the other hand, position queries concentrate on the temporal dimensions of an action. We defer to the supplementary material the detailed encoder architecture.
Query selection. We add a query selection module before passing the encoder output query features to the decoder. We route the query features outputted by the encoder to two respective prediction heads (described later in this section), which independently decode them into predictions for action class and temporal location, as depicted on the right side of Fig. 2. The action head processes the action queries $\tilde{\mathbf{E}}^{\text{act}}$ to estimate the action class $\tilde{\mathbf{A}}$ for each query, while the position head decodes the position queries $\tilde{\mathbf{E}}^{\text{pos}}$ to their temporal locations $\tilde{\mathbf{P}}$. Based on the prediction results, we discard query features classified as “others ($\emptyset$)”, retaining only those identified as “repetitive actions”. For the remaining queries, we rank them based on their classification confidence and preserve only the top $Q$ high-confidence queries. The others are discarded. After this selection process, both the action query features $\mathbf{E}^{\text{act}}$ and position query features $\mathbf{E}^{\text{pos}}$ have a dimension of $Q\times C$. Additional details will be provided when we discuss the prediction head later in this section.
Decoder. The decoder $\mathcal{D}(\cdot)$ is also a classical Transformer [40] architecture, consisting of ${L}^{\text{dec}}$ standard decoder layers. It processes a set of action queries $\mathbf{Q}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and position queries $\mathbf{Q}^{\text{pos}}\in\mathbb{R}^{Q\times C}$, while simultaneously attending to the direct output (i.e. $\tilde{\mathbf{E}}^{\text{act}}$ and $\tilde{\mathbf{E}}^{\text{pos}}$) from the encoder, as depicted in Fig. 2 (right). Following the practice in DETR [3], these inputs are decoded in parallel across each decoder layer and transformed into the corresponding query features $\mathbf{D}^{\text{act}}\in\mathbb{R}^{Q\times C}$ and $\mathbf{D}^{\text{pos}}\in\mathbb{R}^{Q\times C}$. The input position queries $\mathbf{Q}^{\text{pos}}$ are initialized as learnable parameters. To address the open-set challenge in the TRC task, we introduce the DAQ strategy, which initializes the decoder input action queries using the selected encoder embeddings $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$. The details of the DAQ strategy will be elaborated in Sec. 3.4. Using self- and encoder-decoder attention over these embeddings, our method globally reasons selected video frame features together while being able to use the whole frame-wise video feature as context. The two sets of $Q$ queries are then independently transformed into corresponding prediction results, i.e. an action head decodes $\mathbf{D}^{\text{act}}$ into the action class prediction $\mathbf{A}$, and a position head decodes $\mathbf{D}^{\text{pos}}$ into the temporal location prediction $\mathbf{P}$. We detail the two heads in the following subsection.
Prediction heads. The prediction heads are designed to make the final predictions using the features processed by the encoder-decoder. They comprise two networks: an action head and a position head, both of which are Multi-Layer Perceptrons (MLPs).
The action head processes the output action queries from the decoder $\mathbf{D}^{\text{act}}$ to estimate the action class $\mathbf{A}\in\mathbb{R}^{Q\times 1}$ for each query using a softmax function. Each element of $\mathbf{A}$ represents the probability that the corresponding query is a “repetitive action”, with values spanning from 0 to 1. In practice, a threshold value, $\alpha$, is used to categorize the queries: values greater than $\alpha$ classify a query as a “repetitive action” while values less than or equal to $\alpha$ categorize it as “others ($\emptyset$)”. Recall that in the query selection module, we rank the queries based on their estimated probabilities $\tilde{\mathbf{A}}\in\mathbb{R}^{T\times 1}$, which serve as confidence scores. We then select the top $Q$ queries with the highest confidence scores according to the ranking.
The position head estimates the temporal location $\mathbf{P}=(\mathbf{m},\mathbf{d})\in\mathbb{R}^{Q\times 2}$ for each query based on the output position queries from the decoder $\mathbf{D}^{\text{pos}}$. We use the midpoint $\mathbf{m}\in\mathbb{R}^{Q}$ and the duration $\mathbf{d}\in\mathbb{R}^{Q}$ to denote the temporal location of an action cycle.
Bipartite matching. Following the practice of DETR [3], our method produces a fixed-size set of $Q$ predictions in a single pass, where $Q$ is set to be much larger than the typical number of action cycles in a $T$-frame video sequence. Note that the $Q$ predictions are in a random order. One of the main difficulties of training is to match the predicted action cycles (classes and temporal locations) to the GT action cycles, therefore, optimizing action cycle-specific losses, e.g. classification loss, temporal position loss, etc.
Let us denote GT repetitive action cycle labels by $\hat{\mathbf{Y}}=(\hat{\mathbf{A}},\hat{\mathbf{P}})$. $\hat{\mathbf{A}}=\mathbf{1}\in\mathbb{R}^{\hat{N}\times 1}$ denotes the $\hat{N}$ repetitive action class labels, where we assign the value 1 to the “repetitive actions”. $\hat{\mathbf{P}}\in\mathbb{R}^{\hat{N}\times 2}$ denotes the GT temporal positions of the $\hat{N}$ repetitive action, represented by $\hat{N}$ sets of midpoints and duration. Similarly, we denote the predicted repetitive actions by $\mathbf{Y}=(\mathbf{A},\mathbf{P})$, a set of size $Q$ predictions. We set $Q$ much larger than $\hat{N}$ empirically. We consider $\hat{\mathbf{Y}}$ also a set of size $Q$ padded with “others ($\emptyset$)”, i.e. we assign the value 0 to denote $\emptyset$ class. To find a bipartite matching between these two sets $\hat{\mathbf{Y}}$ and $\mathbf{Y}$, we search for a permutation of $Q$ elements $\sigma\in\mathfrak{S}_{Q}$ with the lowest cost:
$\hat{\sigma}=\arg\min_{\sigma\in\mathfrak{S}_{Q}}\sum_{i=1}^{Q}\mathcal{L}_{\text{match}}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{\sigma(i)}),$ (1)
where $\mathcal{L}_{\text{match}}$ is a pair-wise matching cost between GT action cycle $\hat{\mathbf{Y}}_{i}$ and a prediction cycle with index $\sigma(i)$.
The matching cost takes into account both the action classification result and the similarity of the predicted temporal locations and the GT temporal location. Each element $i$ of the GT action cycle set can be seen as $\hat{\mathbf{Y}}_{i}=(\hat{\mathbf{A}}_{i},\hat{\mathbf{P}}_{i})$ where $\hat{\mathbf{A}}_{i}$ is the target class label (1 for “repetitive actions”, and 0 for $\emptyset$) and $\hat{\mathbf{P}}_{i}\in\mathbb{R}^{2}$ is a vector that denotes the midpoint time position and lasting duration of a GT action cycle. For the prediction with index $\sigma(i)$, we define probability of class $\hat{\mathbf{A}}_{i}$ as $p_{\sigma(i)}(\hat{\mathbf{A}}_{i})$. Then we can define the matching cost as
$\begin{split}\mathcal{L}_{\text{match}}(\hat{\mathbf{Y}}_{i},\mathbf{Y}_{\sigma(i)})&=-\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}p_{\sigma(i)}(\hat{\mathbf{A}}_{i})\\ &+\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)}),\end{split}$ (2)
where $\mathbbm{1}$ is an indicator function.
To measure the similarity of the predicted temporal locations and the GT temporal location, we define $\mathcal{L}_{\text{pos}}$ using the linear combination of the $L_{1}$ distance and the generalized Intersection over Union (IoU) loss [32]. Overall, the position loss is defined as
@@
$\begin{split}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)})&=\lambda_{\text{L1}}\|\hat{\mathbf{P}}_{i}-\mathbf{P}_{\sigma(i)}\|_{1}\\ &+\lambda_{\text{gIoU}}\mathcal{L}_{\text{gIoU}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\sigma(i)}),\end{split}$ (3)
where $\lambda_{\text{gIoU}},\lambda_{\text{L1}}\in\mathbb{R}$ are hyperparamters.
By employing the Hungarian matching algorithm [16] to optimize Eq. 1, we can achieve the final optimal matching $\hat{\sigma}$ which uniquely assigns a prediction to a GT action cycle, finding one-to-one matching without duplicates. Notice that the matching cost between a repetitive action instance and $\emptyset$ doesn’t depend on the prediction, which means that in that case the cost is a constant. Following this step, we can apply the corresponding losses related to classification and temporal location prediction.
We define a Hungarian loss for all pairs matched in the previous step to supervise both classification and temporal location predictions, i.e. a linear combination of a negative log-likelihood loss for class prediction and a temporal position loss defined in Eq. 3:
$\begin{split}\mathcal{L}_{\text{Hungarian}}(\hat{\mathbf{Y}},\mathbf{Y})&=\sum\limits_{i=1}^{Q}\left[-\text{log}\,{p_{\hat{\sigma}(i)}}{(\hat{\mathbf{A}}_{i})}\right.\\ &+\left.\mathbbm{1}_{\{\hat{\mathbf{A}}_{i}\neq\emptyset\}}\mathcal{L}_{\text{pos}}(\hat{\mathbf{P}}_{i},\mathbf{P}_{\hat{\sigma}(i)})\right],\end{split}$ (4)
where $\hat{\sigma}$ is the optimal assignment computed in the previous bipartite matching step, i.e. Eq. 1.